OpenStudy (anonymous):

There are seven black socks, eight blue socks, 10 white socks, and five patterned socks in the drawer. What is the probability of pulling a blue sock, replacing it, and then drawing a patterned sock out of the drawer?

OpenStudy (anonymous):

@OregonDuck

oregonduck (oregonduck):

Since there is replacement, the total no. of socks in the drawer does not change, therefore: The probability of pulling a patterned sock on after the replacement= no. of patterned socks/total no. of socks = 5/30 = 1/6

OpenStudy (moazzam07):

7/30 it's 7 black socks

OpenStudy (moazzam07):

Wait no

OpenStudy (anonymous):

im going to start putting the answers up lol 3/30 13/60 2/45 4/9

OpenStudy (moazzam07):

You was right

OpenStudy (moazzam07):

That would be helpful

oregonduck (oregonduck):

well first of all there are 25 socks total/ Considering there are 8 blue socks you have a 8/25 chance of getting a blue sock and because you replace that sock there are still 25 socks and the odds of getting a patterned sock is 5/25. so we just multiply the two together 8/25 * 5/25 = 40/625 simplify to 8/125. I pretty sure that's the answer it's been awhile since I have done this.

OpenStudy (moazzam07):

Your right @OregonDuck

OpenStudy (anonymous):

thats not a choice tho

OpenStudy (anonymous):

@heretohelpalways

OpenStudy (anonymous):

I am here

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

yes, what is the probability of each event happening by itself?

OpenStudy (anonymous):

1/30

OpenStudy (anonymous):

and what are the options?

OpenStudy (moazzam07):

Ok it's 8 blue socks and 5 pattern socks

OpenStudy (anonymous):

3/30 16/60 2/45 4/9

OpenStudy (anonymous):

so if it is 8/30 probability for blue socks and 5/30 for patterned socks, multiply 8/30 by 5/30

OpenStudy (moazzam07):

The probability for picking blue and replacing it would be 13/30

OpenStudy (anonymous):

? that would not work im so confused

OpenStudy (moazzam07):

30 divided by 13 is 2.4

OpenStudy (anonymous):

still no :(

OpenStudy (moazzam07):

What are the options

OpenStudy (anonymous):

ahhhhhhhhhhhh

OpenStudy (anonymous):

so the answer would be 2/45

OpenStudy (anonymous):

why would it be 2/45

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

because 8/30*5/30=2/45

OpenStudy (moazzam07):

Yes heretohelpalways it right

OpenStudy (moazzam07):

C is the correct answer

OpenStudy (anonymous):

to find the probability of a compound event, multiply each independent event

OpenStudy (anonymous):

in this case, the independent event probabilities are 8/30 and 5/30